# Graph-based Square-Root Estimation for Sparse Linear Regression

### v1

@@ 2.3 Choice of the tuning parameter
For a given constant $r>1$, we set $\bar{r}:=\frac{r+1}{r-1}$. For given $\lambda>0$, we define the set
${\mathcal{A}}:=\bigl{\{}V\leq\lambda/\bar{r}\bigr{\}}.$ (2.4)
Next, we will prove that, for any $\alpha$ close to zero and a suitable tuning parameter $\lambda$, the set ${\mathcal{A}}$ holds with probability $1-\alpha$. Before presenting the specific conclusion, we first provide the following assumption:
@@
$\lambda\geq\frac{n\bar{r}\sqrt{2\zeta d_{max}}}{\tau_{min}\sqrt{n-d_{max}}}\Big{(}1+\sqrt{\frac{2log(2p/\alpha)}{d_{min}}}\Big{)},$ (2.5)
@@
It can be noted that the analysis in this part only considers Gaussian noise, but the results can also be generalized to different types of noise by applying different deviation inequalities [6, 27, 42]. For example, if $\epsilon_{i}$’s belong to a general sub-exponential family, the order of magnitude of $\lambda$ remains the same. For more analysis on non-Gaussian noise, one can refer to [3]. A detailed study is omitted here.
@@ 2.4 Finite sample bounds
In this subsection, we will give the results of finite sample bounds for the proposed GSRE method over the set ${\mathcal{A}}$. The following assumptions need to be considered in subsequent analysis.
@@
Assumption 2.2 guarantees that predictors connected to the useful predictors are also useful, which is reasonable. Assumption 2.3 is a general compatibility condition on the design matrix [7, 8] and is a slight relaxation of the widely used restricted eigenvalue condition [4]. The constant $\kappa$ quantifies the correlations in the design matrix: a smaller value of $\kappa$ indicates stronger correlations. Assumption 2.4 essentially constrains the true model to have fewer parameters than $n$.
@@
Assumptions 2.1, 2.2, 2.3, 2.4 hold. Then, on the event ${\mathcal{A}}$, for any optimal solution $\hat{\beta}$ of model (1.4), we have
@@
Theorem 2.1 conducts an estimation analysis based on a tuning parameter independent of $\sigma$ and yields results similar to existing literature related to group lasso and group square-root lasso estimators [31, 8]. Furthermore, it is noted that, from Theorem 2.1, subset recovery can be guaranteed with no noise, i.e., $\sigma=0$. However, for classical least squares methods such as lasso, elastic net and group lasso, even if $\sigma=0$, accurate recovery cannot be guaranteed because their tuning parameters are heavily dependent on $\sigma$. Additionally, combining with Lemma 2.3 on the selection of regularization parameter, it can be concluded that the result of Theorem 2.1 holds with probability $1-\alpha-(1-\alpha)(1+e^{2})e^{-n/24}$.
@@ 2.5 Asymptotic normality and model selection consistency
The errors $\epsilon_{1}$, $\epsilon_{2}$, $\ldots$, $\epsilon_{n}$ are i.i.d. variables with mean 0 and variance $1$.
@@
Theorem 2.2 also provides a guideline on how to choose the positive weight $\tau_{i}$. [34] suggest to consider $\tau_{i}=d_{i}^{\eta}$ for some $\eta\in\big{(}0,\frac{1}{2}\big{)}$. More specially, they give a critical value with $\eta=\frac{log(2)}{2log(3)}$, which is the smallest value that it becomes impossible to select a support of exactly two covariates. In our numerical experiments, we also choose $\tau_{i}=d_{i}^{\eta}$ with $\eta=\frac{log(2)}{2log(3)}$.
@@
Assumption 2.7 ensures the sparsity of the problem. Assumption 2.8 can be achieved by normalizing each predictor. Assumption 2.9 guarantees that matrix $\frac{X_{{\mathcal{I}}_{*}}^{\top}X_{{\mathcal{I}}_{*}}}{n}$ is invertible. Assumption 2.10 is similar to the strong irrepresentable condition [55].
@@
Then, on the set ${\mathcal{A}}$, as $n\to\infty$ and $p=p(n)\to\infty$, there exists a solution $\hat{\beta}$ to (1.4) such that $\operatorname*{sign}(\hat{\beta})=\operatorname*{sign}(\beta^{*})$ with probability tending to 1.
Proof.
See Appendix F. ∎
@@ 3 Computation
Upon careful observation, we note that model (1.4) has two non-smooth convex structures. It is well known that ADMM is a flexible and efficient optimization method that can handle convex optimization problems with separable structures and linear constraints. Therefore, we use ADMM for the computation here. First, to effectively utilize ADMM, we introduce two auxiliary variables to reformulate model (1.4) as the following minimization problem with two-block separable structures and two linear constraints:
$\begin{array}[]{lll}\min\limits_{\beta,u,v}&\frac{\|u\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|v\|_{G,\tau}\\ \text{s.t.}&X\beta-u=y,\\ &\beta-v=0.\end{array}$ (3.1)
Given $\sigma>0$, the augmented Lagrangian function of (3.1) is defined by
@@
In order to avoid the time-consuming issue of inverting large matrices in high-dimensional environments, we use the Sherman-Morrison-Woodbury formula [20] in specific computations, i.e., $\Big{(}{\mathcal{I}}_{p}+X^{\top}X\Big{)}^{-1}={\mathcal{I}}_{p}-X^{\top}({\mathcal{I}}_{n}+XX^{\top})^{-1}X$. This process successfully transfers the inverse operation of a $p\times p$ matrix to a $n\times n$ matrix, which greatly reduces the computational cost of the algorithm when $p$ is much larger than $n$.
@@
Denote $r^{k+1}:=X\beta^{k+1}-y+\sigma^{-1}z^{k}$. From [1, Example 6.19], we have
@@
where ${\mathcal{O}}^{k+1}:=\left\{i\in[p]\ |\ \|h^{k+1}_{{\mathcal{N}}_{i}}\|_{2}>\frac{\lambda\tau_{i}}{\sigma n}\right\}$, $S_{{\mathcal{O}}^{k+1}}:=\left\{x\in\mathbb{R}^{p}\ |\ \|x_{{\mathcal{N}}_{i}}\|_{2}\leq\frac{\lambda\tau_{i}}{\sigma n}\text{ for each }i\in{\mathcal{O}}^{k+1}\right\}$. The main problem we face is to solve the projection of $h^{k+1}$ onto the convex set $S_{\mathcal{O}^{k+1}}$. Similar to [50], we also use different methods to calculate flexibly based on the number of elements in ${\mathcal{O}}^{k+1}$, denoted as $|{\mathcal{O}}^{k+1}|$. Specifically, when $|{\mathcal{O}}^{k+1}|<p/10$, we use the Bertseka’s projected Newton method [44] to calculate this projection from the dual perspective. However, when $|{\mathcal{O}}^{k+1}|\geq p/10$, we use the Parallel Dykstra-like proximal algorithm [10]. The details about these two algorithms are shown in the supplementary materials of [50].
Based on the above analysis, we see that each minimization subproblem has been effectively computed, and thus we can summarize the algorithm framework of ADMM for model (3.1).
@@
The non-interdependence of the $u$ and $v$ subproblems makes the above iterative framework equivalent to 2-block ADMM with guaranteed convergence. Specific convergence results are detailed in [15], and will not be further elaborated here.
@@ 4 Numerical experiments
To thoroughly evaluate the numerical performance of the proposed GSRE method, we compare it with several classic linear regression methods. Specifically, we compare it with well-established ridge regression (Ridge), lasso, adaptive lasso (Alasso) and elastic net (Enet), which do not utilize the structural information of the predictor graph. Their results are obtained by running their corresponding internal functions in MATLAB. Additionally, to highlight the advantages of the graph penalty in this paper, we compare GSRE with the traditional square-root lasso (SRL) in [3]. Furthermore, to demonstrate the advantages of the square-root loss in this paper, we compare GSRE with the graph sparse regression based on least squares loss, i.e., SRIG in [50].
@@ 4.1 Simulation data
In the simulation studies, the predictor graph is defined by the precision matrix of the predictors. In practice, the predictor graph is estimated by the graphical lasso method [18]. We also consider the performance of GSRE and SRIG with the true predictor graph, denoted as GSRE-o and SRIG-o, respectively. In the entire simulation experiment, we uniformly set $p=100$ and the true coefficient vector as $\beta^{*}=(3,\ldots,3,0,\ldots,0)^{\top}$ and $s^{*}=15$. The simulated data consists of a training set, an independent validation set, and an independent test set. All models are fitted only on the training data. The validation set is used to choose the appropriate tuning parameter, and we use the BIC criterion to determine the optimal tuning parameter. The test set is used to evaluate the performance of different methods. We use the notation ./././ to denote the sample sizes in the training, validation and test sets, respectively. To highlight $p>n$, for each example, we consider three cases: (I) 20/20/400, (II) 40/40/400 and (III) 60/60/400. For each case, we repeat the simulation 50 times to observe the stability of the algorithms.
@@
Figure 1 illustrates the true predictor graphs of the three simulated examples. Since Examples 1 and 2 have the same correlation among predictors except for differences in their internal distributions, their predictor graph structures are identical. Since there are three groups of correlations among the first 15 predictors, the predictor graph contains three interconnected subgraphs. Because the remaining predictors are independent of each other, the predictor graph includes 85 isolated nodes. Due to their similarity, we have combined their experimental results. Assuming the precision matrix corresponding to the predictor graph is denoted by $\Omega=(\omega_{ij})_{i,j=1,2,\ldots,p}$, the generation method for Example 3 leads to $\omega_{ii}=1.333$, $\omega_{ij}=-0.667$ if $|i-j|=1$ and $\omega_{ij}=0$ if $|i-j|>1$. This indicates that predictors are only correlated with variables in their neighborhood, resulting in a banded structure for the predictor graph.
@@
  • Normal distribution: $\epsilon\sim N(0,5^{2})$.
  • T-distribution: $\epsilon\sim t(10)/2\sqrt{5}$.
@@
Here, we consider normal distribution with high noise levels, t-distribution with heavier tails, laplace distribution with high kurtosis, and uniform distribution presenting a scenario with constant probability. These all create challenging statistical environments. To evaluate the performances of different methods, we use the following measures:
@@
  • False positive rate (FPR) and false negative rate (FNR).
@@
Tables 1-4 show the performance of each method for Examples 1 and 2 under four different types of noise. The numbers in parentheses represent the corresponding standard deviations, which are used to measure the stability of these methods. A review of the results in the tables shows that the patterns presented in the four tables are similar. Specifically, among the five methods that do not utilize the predictor graph structure information, Enet achieves better estimation results. Alasso performs better in both RPE and FPR. Although Ridge has the smallest FNR, its FPR is nearly 1. This phenomenon indicates that nearly all zero elements are estimated as non-zero elements. In cases (I) and (II), where $p\gg n$, the proposed GSRE method generally performs best in estimation, prediction and model selection for Example 1 and Example 2. Although in case (I), the FNR of GSRE is slightly higher than that of SRIG for Example 2, it is noteworthy that our FPR is significantly lower than that of SRIG. This observation clearly indicates that our method also outperforms SRIG in model selection. For case (III), where the sample size is slightly larger, GSRE remains optimal in model estimation and prediction, but is slightly inferior to SRIG in model selection. Additionally, we find that for Example 1 and Example 2, the results using the estimated predictor graph and the true predictor graph are almost identical, and sometimes the performance with the true predictor graph is slightly worse than that with the estimated predictor graph. This discrepancy may be due to inherent precision issues in MATLAB when computing the inverse of the covariance matrix corresponding to the true predictor graph.
Table 1: Performance comparison of estimation, prediction and model selection for Example 1&2 with normal distribution noise.
Example 1 Example 2
$L_{2}$ distance RPE FPR FNR $L_{2}$ distance RPE FPR FNR
(I) 20/20/400 Ridge 8.177(0.726) 7.130(2.433) 0.966(0.016) 0.000(0.000) 8.848(0.617) 7.700(1.952) 0.967(0.013) 0.001(0.009)
Lasso 12.899(1.820) 4.380(3.754) 0.097(0.023) 0.511(0.120) 11.390(1.230) 6.241(3.121) 0.092(0.020) 0.507(0.117)
Alasso 12.504(1.616) 3.253(2.356) 0.081(0.021) 0.441(0.105) 11.246(1.326) 5.248(2.820) 0.080(0.019) 0.440(0.104)
Enet 6.930(1.193) 5.419(2.932) 0.435(0.050) 0.011(0.034) 8.119(1.046) 6.954(2.435) 0.430(0.049) 0.036(0.069)
SRL 12.558(1.768) 3.842(2.837) 0.103(0.028) 0.464(0.110) 11.172(1.524) 5.836(3.588) 0.099(0.029) 0.439(0.118)
SRIG-o 8.773(0.656) 13.543(2.324) 0.565(0.063) 0.000(0.000) 9.305(0.513) 12.236(1.691) 0.565(0.058) 0.007(0.031)
SRIG 8.818(0.647) 13.702(2.324) 0.567(0.061) 0.000(0.000) 9.235(0.518) 12.060(1.691) 0.564(0.056) 0.009(0.049)
GSRE-o 3.638(1.470) 2.110(3.027) 0.022(0.020) 0.000(0.000) 4.817(1.513) 2.856(3.030) 0.023(0.022) 0.013(0.066)
GSRE 3.426(1.301) 1.888(2.525) 0.025(0.023) 0.000(0.000) 4.775(1.531) 2.856(3.056) 0.023(0.021) 0.013(0.066)
(II) 40/40/400 Ridge 6.972(0.614) 3.640(1.045) 0.960(0.015) 0.000(0.000) 7.334(0.581) 3.798(0.790) 0.966(0.013) 0.000(0.000)
Lasso 9.181(1.344) 1.567(0.699) 0.056(0.018) 0.251(0.091) 7.722(1.277) 2.359(1.100) 0.045(0.019) 0.192(0.104)
Alasso 9.448(1.388) 1.030(0.434) 0.049(0.018) 0.245(0.090) 7.771(1.337) 1.573(0.712) 0.040(0.018) 0.180(0.091)
Enet 5.330(0.692) 4.007(1.321) 0.310(0.054) 0.000(0.000) 6.008(0.744) 4.223(1.269) 0.316(0.045) 0.005(0.018)
SRL 9.098(1.343) 1.455(0.646) 0.067(0.025) 0.243(0.085) 7.607(1.283) 2.274(1.042) 0.051(0.027) 0.180(0.096)
SRIG-o 3.312(1.136) 1.778(1.779) 0.031(0.058) 0.000(0.000) 4.120(1.335) 2.325(2.080) 0.054(0.076) 0.000(0.000)
SRIG 3.118(1.180) 1.737(1.726) 0.031(0.059) 0.000(0.000) 3.860(1.184) 1.999(1.724) 0.044(0.065) 0.000(0.000)
GSRE-o 2.487(0.505) 0.667(0.439) 0.012(0.013) 0.000(0.000) 3.041(0.594) 0.783(0.505) 0.016(0.013) 0.000(0.000)
GSRE 2.385(0.529) 0.659(0.446) 0.013(0.013) 0.000(0.000) 3.007(0.576) 0.775(0.464) 0.016(0.013) 0.000(0.000)
(III) 60/60/400 Ridge 6.174(0.738) 2.221(0.579) 0.961(0.012) 0.000(0.000) 6.280(0.744) 2.235(0.551) 0.958(0.012) 0.000(0.000)
Lasso 7.541(1.227) 0.940(0.335) 0.036(0.016) 0.125(0.070) 5.786(1.224) 1.248(0.669) 0.024(0.015) 0.051(0.058)
Alasso 8.708(1.405) 0.682(0.232) 0.036(0.015) 0.171(0.071) 6.504(1.152) 0.960(0.365) 0.022(0.012) 0.085(0.063)
Enet 4.739(0.674) 3.559(1.181) 0.204(0.056) 0.000(0.000) 5.358(0.894) 3.756(1.477) 0.198(0.043) 0.001(0.009)
SRL 7.423(1.205) 0.897(0.362) 0.039(0.019) 0.120(0.066) 5.647(1.163) 1.174(0.579) 0.029(0.020) 0.040(0.057)
SRIG-o 2.489(0.426) 0.761(0.374) 0.005(0.008) 0.000(0.000) 3.245(0.700) 1.303(0.703) 0.001(0.004) 0.000(0.000)
SRIG 2.259(0.446) 0.721(0.383) 0.005(0.008) 0.000(0.000) 2.656(0.562) 0.743(0.430) 0.004(0.006) 0.000(0.000)
GSRE-o 2.109(0.313) 0.381(0.188) 0.008(0.011) 0.000(0.000) 2.453(0.442) 0.452(0.229) 0.008(0.010) 0.000(0.000)
GSRE 2.016(0.317) 0.377(0.185) 0.009(0.010) 0.000(0.000) 2.420(0.448) 0.443(0.224) 0.008(0.010) 0.000(0.000)

### v2

@@ 2.3 Choice of the tuning parameter
For a given constant $r>1$, we set $\bar{r}:=\frac{r+1}{r-1}$. Given $\lambda>0$, we define
${\mathcal{A}}:=\bigl{\{}V\leq\lambda/\bar{r}\bigr{\}},$ (8)
which represents a event that inequality $V\leq\lambda/\bar{r}$ holds. Next, we will prove that, for any $\alpha$ close to zero and a suitable tuning parameter $\lambda$, the event ${\mathcal{A}}$ holds with probability $1-\alpha$. Before presenting the specific conclusion, we first provide the following assumption:
@@
$\lambda\geq\frac{n\bar{r}\sqrt{2\zeta d_{max}}}{\tau_{min}\sqrt{n-d_{max}}}\Big{(}1+\sqrt{\frac{2log(2p/\alpha)}{d_{min}}}\Big{)},$ (9)
@@
It can be noted that the analysis in this part only considers Gaussian noise, but the results can also be generalized to different types of noise by applying different deviation inequalities (Bousquet, 2002; Lederer and van de Geer, 2011; van de Geer and Lederer, 2013). For example, if $\epsilon_{i}$’s belong to a general sub-exponential family, the order of magnitude of $\lambda$ remains the same. The corresponding analysis can be found in the supplementary materials. For more analysis on non-Gaussian noise, one can refer to Belloni et al. (2011). A detailed study is omitted here.
@@ 2.4 Finite sample bounds
In this subsection, we will give the results of finite sample bounds for the proposed GSRE method over the event ${\mathcal{A}}$. The following assumptions need to be considered in subsequent analysis.
@@
Assumption 2.2 guarantees that predictors connected to the useful predictors are also useful, which is reasonable. Assumption 2.3 is a general compatibility condition on the design matrix (Bühlmann and Geer, 2011; Bunea et al., 2013) and is a slight relaxation of the widely used restricted eigenvalue condition (Bickel et al., 2009). The constant $\kappa$ quantifies the correlations in the design matrix: a smaller value of $\kappa$ indicates stronger correlations. An example where Assumption 2.3 holds is presented in the supplementary materials. Assumption 2.4 essentially constrains the true model to have fewer parameters than $n$.
@@
Assumptions 2.1, 2.2, 2.3, 2.4 hold. Then, on the event ${\mathcal{A}}$, for any optimal solution $\hat{\beta}$ of model (4), we have
@@
Theorem 2.1 gives similar results to group lasso and group square-root lasso estimators (Lounici et al., 2011; Bunea et al., 2013). It is worth noting that the estimation analysis is based on a $\sigma$-free tuning parameter, which is obviously different from SRIG in Yu and Liu (2016). Furthermore, it is noted that, from Theorem 2.1, subset recovery can be guaranteed with no noise, i.e., $\sigma=0$. However, for classical least squares methods such as lasso, elastic net and group lasso, even if $\sigma=0$, accurate recovery cannot be guaranteed because their tuning parameters are heavily dependent on $\sigma$. In addition, an empirical verification of $\|\hat{\beta}-\beta^{*}\|_{2}$ can be found in the supplementary materials. Additionally, combining with Lemma 2.3 on the selection of regularization parameter, it can be concluded that the result of Theorem 2.1 holds with probability $1-\alpha-(1-\alpha)(1+e^{2})e^{-n/24}$.
@@ 2.5 Asymptotic normality and model selection consistency
The errors $\epsilon_{1}$, $\epsilon_{2}$, $\ldots$, $\epsilon_{n}$ are i.i.d. variables with mean 0 and variance 111.
@@
Theorem 2.2 also provides a guideline on how to choose the positive weight $\tau_{i}$. Obozinski et al. (2011) suggest to consider $\tau_{i}=d_{i}^{\eta}$ for some $\eta\in\big{(}0,\frac{1}{2}\big{)}$. More specially, they give a critical value with $\eta=\frac{log(2)}{2log(3)}$, which is the smallest value that it becomes impossible to select a support of exactly two covariates. In our numerical experiments, we also choose $\tau_{i}=d_{i}^{\eta}$ with $\eta=\frac{log(2)}{2log(3)}$.
@@
Assumption 2.7 ensures the sparsity of the problem. Assumption 2.8 can be achieved by normalizing each predictor. Assumption 2.9 guarantees that matrix $\frac{X_{{\mathcal{I}}_{*}}^{\top}X_{{\mathcal{I}}_{*}}}{n}$ is invertible. Assumption 2.10 is similar to the strong irrepresentable condition (Zhao and Yu, 2006).
@@
Then, on the event ${\mathcal{A}}$, as $n\to\infty$ and $p=p(n)\to\infty$, there exists a solution $\hat{\beta}$ to (4) such that $\operatorname*{sign}(\hat{\beta})=\operatorname*{sign}(\beta^{*})$ with probability tending to 1.
@@ 3 Computation
Upon careful observation, we note that model (4) has two non-smooth convex structures. It is well known that ADMM is a flexible and efficient optimization method that can handle convex optimization problems with separable structures and linear constraints. Therefore, we use ADMM for the computation here. First, to effectively utilize ADMM, we introduce two auxiliary variables to reformulate model (4) as the following minimization problem with two-block separable structures and two linear constraints:
$\begin{array}[]{lll}\min\limits_{\beta,u,v}&\frac{\|u\|_{2}}{\sqrt{n}}+\frac{\lambda}{n}\|v\|_{G,\tau}\\ \text{s.t.}&X\beta-u=y,\\ &\beta-v=0.\end{array}$ (11)
Given $\sigma>0$, the augmented Lagrangian function of (11) is defined by
@@
In order to avoid the time-consuming issue of inverting large matrices in high-dimensional environments, we use the Sherman-Morrison-Woodbury formula (Golub and Loan, 2013) in specific computations, i.e., $\Big{(}{\mathcal{I}}_{p}+X^{\top}X\Big{)}^{-1}={\mathcal{I}}_{p}-X^{\top}({\mathcal{I}}_{n}+XX^{\top})^{-1}X$. This process successfully transfers the inverse operation of a $p\times p$ matrix to a $n\times n$ matrix, which greatly reduces the computational cost of the algorithm when $p$ is much larger than $n$.
@@
Denote $r^{k+1}:=X\beta^{k+1}-y+\sigma^{-1}z^{k}$. From Beck (2017, Example 6.19), we have
@@
where ${\mathcal{O}}^{k+1}:=\left\{i\in[p]\ |\ \|h^{k+1}_{{\mathcal{N}}_{i}}\|_{2}>\frac{\lambda\tau_{i}}{\sigma n}\right\}$, $S_{{\mathcal{O}}^{k+1}}:=\left\{x\in\mathbb{R}^{p}\ |\ \|x_{{\mathcal{N}}_{i}}\|_{2}\leq\frac{\lambda\tau_{i}}{\sigma n}\text{ for each }i\in{\mathcal{O}}^{k+1}\right\}$. The main problem we face is to solve the projection of $h^{k+1}$ onto the convex set $S_{\mathcal{O}^{k+1}}$. Similar to Yu and Liu (2016), we also use different methods to calculate flexibly based on the number of elements in ${\mathcal{O}}^{k+1}$, denoted as $|{\mathcal{O}}^{k+1}|$. Specifically, when $|{\mathcal{O}}^{k+1}|<p/10$, we use the Bertseka’s projected Newton method (Villa et al., 2014) to calculate this projection from the dual perspective. However, when $|{\mathcal{O}}^{k+1}|\geq p/10$, we use the Parallel Dykstra-like proximal algorithm (Combettes and Pesquet, 2011). The details about these two algorithms are shown in the supplementary materials of Yu and Liu (2016).
Based on the above analysis, we see that each minimization subproblem has been effectively computed, and thus we can summarize the algorithm framework of ADMM for model (11).
@@
The non-interdependence of the $u$ and $v$ subproblems makes the above iterative framework equivalent to 2-block ADMM with guaranteed convergence. Specific convergence results are detailed in Fazel et al. (2013), and will not be further elaborated here.
@@ 4 Numerical experiments
To thoroughly evaluate the numerical performance of the proposed GSRE method, we compare it with several classic linear regression methods. Specifically, we compare it with well-established ridge regression (Ridge), lasso, adaptive lasso (Alasso) and elastic net (Enet), which do not utilize the structural information of the predictor graph. Their results are obtained by running their corresponding internal functions in MATLAB. Additionally, to highlight the advantages of the graph penalty in this paper, we compare GSRE with the traditional square-root lasso (SRL) in Belloni et al. (2011). Furthermore, to demonstrate the advantages of the square-root loss in this paper, we compare GSRE with the graph sparse regression based on least squares loss, i.e., SRIG in Yu and Liu (2016).
@@ 4.1 Simulation data
In the simulation studies, the predictor graph is defined by the precision matrix of the predictors. Specifically, we employ a MATLAB implementation graphicalLasso.m, which uses a graphical Lasso method (Friedman et al., 2008) for estimating sparse precision matrix and can be available at https://github.com/xiaohuichen88/Graphical-Lasso/tree/main. It is worth noting that the precision matrix has already been symmetrized by the previous procedure. In addition, the supplementary materials provide a sensitivity analysis of GSRE to inaccuracies in the estimated graph. We also consider the performance of GSRE and SRIG with the true predictor graph, denoted as GSRE-o and SRIG-o, respectively. In the entire simulation experiment, we uniformly set $p=100$ and the true coefficient vector as $\beta^{*}=(3,\ldots,3,0,\ldots,0)^{\top}$ and $s^{*}=15$. For the sensitive parameter $\lambda$, we adopt the grid search method. Solution paths are computed for $\lambda/(\sqrt{n/2}\|X\|)=2^{-13},2^{-12.8},\ldots,2^{-5.2},2^{-5}$, which refers to Bunea et al. (2013). This grid is empirically constructed to cover all potentially interesting solutions. Then we use the HBIC criterion (Wang et al., 2013) to select the optimal tuning parameter. The simulated data consists of a training set, an independent validation set, and an independent test set. All models are fitted only on the training data. The validation set is used to choose the appropriate tuning parameter. The test set is used to evaluate the performance of different methods. We use the notation ./././ to denote the sample sizes in the training, validation and test sets, respectively. To highlight $p>n$, for each example, we consider three cases: (I) 20/20/400, (II) 40/40/400 and (III) 60/60/400. For each case, we repeat the simulation 50 times to observe the stability of the algorithms.
@@
[…]
@@
  • Gaussian distribution: $\epsilon\sim N(0,5^{2})$.
  • T-distribution: $\epsilon\sim t(2)$.
@@
Here, we consider Gaussian distribution with high noise levels, t-distribution with heavier tails, laplace distribution with high kurtosis, and uniform distribution presenting a scenario with constant probability. These all create challenging statistical environments. To evaluate the performances of different methods, we use the following measures:
@@
  • False positive rate (FPR)= $\frac{FP}{FP+TN}$, false negative rate (FNR)= $\frac{FN}{TP+FN}$ and Mattews correlation coefficient (MCC)= $\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$, where $TP$, $TN$, $FP$, and $FN$ denote the number of true positives, true negatives, false positives, and false negatives, respectively.
@@
Tables 4.1-4 show the performance of each method for Examples 1 and 2 under four different types of noise. The numbers in parentheses represent the corresponding standard deviations, which are used to measure the stability of these methods. A review of the results in the tables shows that the patterns presented in the four tables are similar. Specifically, among the five methods that do not utilize the predictor graph structure, Enet achieves better estimation results. Alasso performs better in RPE. Although Ridge has the smallest FNR, its FPR is nearly 1. This phenomenon indicates that nearly all zero elements are estimated as non-zero elements. In addition, we find that the results obtained by using the estimated predictor graph are the same as those obtained by using the true predictor graph, which indicates that we have achieved accurate estimation of the graph structure between predictors in these simulation examples. In cases (I) and (II), where $p\gg n$, the proposed GSRE method generally performs best in estimation, prediction and model selection for Example 1 and Example 2. For case (III), where the sample size is slightly larger, GSRE remains optimal in estimation and prediction, but is slightly inferior to SRIG in model selection.
Tables 1-4 display the performance of each method for Example 3 under four different types of noise. Similarly, as the sample size increases, the numerical performance of each algorithm improves. Ridge often gives a low FNR but a high FPR, which indicate a tendency towards overfitting. Lasso and Alasso methods generally have lower values of FPR, but they have higher values for $L_{2}$ distance and FNR. Enet still performs best in the methods that don’t use the graph structure between predictors. In all four challenging environments, the GSRE method consistently outperforms other methods in model estimation and prediction, which shows its robust performance. By balancing the values of FPR, FNR and MCC, it is clear that our method also performs very well in model selection for Example 3.